OpenStudy (megannicole51):

yankeecm1005 Randall has an account balance of $675.54 in a savings account that earns interest at a rate of 2.5% compounded twice a year. If Randall opened the account eleven years ago, what was the principal deposit rounded to the nearest dollar?

OpenStudy (anonymous):

\[A = P(1 = \frac{ r }{ n })^{nt}\] \[675.54 = P(1 + \frac{ 0.025 }{ 2 })^{2(11)}\] Just solve for P and you have your answer!

OpenStudy (anonymous):

Thank you for letting me help you! Signed: \[Mangorox\]